GKX Industries has introduced a new family of products to the market, hydraulic seals (in inch and metric sizes). The company expects sales of these products to increase according to the cash flow sequence $80,000 + $6,000 k, where k is in years, starting in year k=1 If i =11% per year. a) What is the future worth of the entire cash flow series 12 years from now? b) What is the equivalent Uniform annual worth (A) for the first 6 years?
Answer #1

The cash flow occurs in a series given by 80000 in all years and 6000K for K years.

a) Future worth (K = 12) = (80000(P/A, 11%, 12) + 6000(P/G, 11%, 12))(F/P, 11%, 12)

= 80000(F/A, 11%, 12) + 6000(P/G, 11%, 12)(F/P, 11%, 12)

= 80000*22.7132 + 6000*27.8388*3.4985

= $2,817,566

b) Equal annual worth for first 6 years = (80000(P/A, 11%, 6) + 6000(P/G, 11%, 6))(A/P, 11%, 6)

= 80000 + 6000(A/G, 11%, 6)

= 80000 + 6000*2.1976

= $93,185.60